# Diffusive Spreading Across Dynamic Mitochondrial Network Architectures

### v1

@@ Results
where the first term in the updated $S$ accounts for material in the source cluster and the second term represents material in the rest of the network. Note that $h_{c}$ now depends on $h_{0}^{(d)},f$. When clusters are large and interactions are infrequent, the above solution reduces to the static network limit of Eq. 1. When the clusters are small, the source cluster is fully filled ($h_{0}^{(d)}\rightarrow h_{0}$) and so is each cluster that interacts with it ($f\rightarrow 1$). The general solution then approaches the fragmented social network limit of Eq. 3.
In Fig. 4, we show the spreading of material through simulated networks with increasing connectivity, spanning across the different regimes. In the limit of rapid decay (small $\tau_{\textrm{d}}$), material spreads within a single cluster of dimension $d$, with the amount delivered scaling as $S\sim\tau_{d}^{d/2}$ When particle diffusivity is fast and clusters are small, the source cluster fills faster than interactions can occur and the amount of material plateaus in the regime of $\tau_{c}<\tau_{d}<\tau_{\text{int}}$. For longer decay times, interaction between distinct clusters dominates the spread, and the total material scales linearly as $S\sim\tau_{d}$. When particle diffusion is slow (Fig. 4b), the cluster containing the source unit is able to offload material through interactions even before it is fully filled ($\tau_{\text{int}}<\tau_{c}$), and the plateau region narrows or disappears.
The simplified analytic model (Eq. 4) approximately matches simulations over a broad range of network connectivities and timescales (Fig. 4). However, some discrepancy occurs for networks that are close to the percolation transition (light green curve), particularly when the particle diffusivity is low. In this case, the approximation systematically overestimates material spreading through the network. This may be the result of transient fissions within a single cluster (which could temporarily hinder material spreading [22]), or steric inaccessibility of nodes buried within a cluster, neither of which are accounted for in our analytic models. Additionally, the percolation transition corresponds to a broad variability in cluster sizes, resulting in the source unit occasionally being trapped in very small clusters that limit material delivery.
@@ Discussion
The above calculations yield succinct, conceptually interpretable predictions for the rate of material spread in dynamic networks. We examine two limiting regimes: one of static, well-connected physical networks and one of socially-interacting homogeneous clusters. In physical networks, spreading increases with material diffusivity and network dimensionality. In social networks, material accumulates linearly in time and is limited either by the fusion rate between nearby clusters or the mobility of those clusters.
@@
The models presented here are based on measurable structural and dynamic features. The static model requires network dimensionality, cluster size, and particle diffusivity, while the social model additionally needs the fusion rate and cluster mobility. Modern imaging techniques allow quantification of mitochondrial network structure and dynamics in a variety of cellular systems  [72, 42, 73, 14, 29]. Our results connect these morphological measurements to the dynamics of material spreading through the mitochondrial population, which is experimentally more challenging to assess. The scaling relationships could in principle be tested by quantifying the dispersion of locally photoconverted proteins through mitochondrial networks of different architectures [35].
This work brings us closer to understanding the functional implications of variable network connectivity for the spread of ions, lipids, proteins, and genetic information within mitochondria. Mammalian cells provide ample examples of mitochondrial network structures in different regimes, from fragmented networks in cells grown on excess glucose [35] or endowed with a Down’s syndrome genotype [34], to networks balanced on the cusp of percolation [14, 73, 58], to hyperfused networks observed in starved cells [35]. A quantitative framework for material spreading in different structures provides insight on the implications of these transitions for many biologically critical processes, such as the distribution of harmful reactive oxygen species [48], the ability to isolate or complement deleterious DNA mutations [74], and the tunneling of calcium ions, proton gradients, and ATP throughout the cell [75, 49, 50].
@@
Our quantitative calculations focus on a simplified dynamic system, with a single source of material held at fixed concentration. Realistic biological scenarios are of course likely to involve additional complications, including interfering sources, non-trivial regulation at the source, or diffusive barriers within the mitochondria [62]. We also assumed that individual mitochondrial units are all identical in their fusion and fission behavior, whereas it is possible that some units remain isolated from the rest of the network, while others are more likely to engage in interactions. However, the minimal model presented here provides a basic building block from which more complicated scenarios can be constructed and which can serve as a null hypothesis for analyzing experimental observations.
The model makes general testable predictions that may be compared against future measurements. We predict that in a fragmented social network, spreading rates should be approximately the same for particles of different diffusivity. In highly-connected physical networks, branched architectures should give rise to more rapid material spread, with a steeper scaling in time, than linear networks. The models presented here could be extended to make quantitative predictions of network filling from multiple source points. For example, this approach could be used to explore how the mitochondrial network fills with calcium that enters at many localized contact sites with the endoplasmic reticulum [76]. Similarly, it enables predictions of how the distribution of mRNA localization and protein import sites on the mitochondrial surface [77, 78] translates into heterogeneity in protein levels within the mitochondria.
@@
While we have focused on mitochondria, analogous problems of diffusive spreading arise for porous media in the geosciences [24], trophallactic networks in honeybees [4], and Alzheimers progression across connected regions in the brain [79]. By bringing together models of static physical networks and transiently interacting social networks, our results enable a quantitatively predictive link from network architecture to the rate of diffusive dispersion.
@@ Acknowledgements.
Funding for this work was provided by NSF grant PHYS-2310229, by the Chan Zuckerberg Initiative, and by a grant from the UCSD Academic Senate. We are grateful to Nigel Goldenfeld and to the members of the CZI Theory Institute Without Walls for helpful discussions.
@@
$\begin{split}S=\frac{\lambda_{p}}{\ell_{0}}\left[\tanh(R_{n}/2\lambda_{p})+\tanh(3R_{n}/2\lambda_{p})\right]\end{split}$ (7)
@@
An additional adjustment accounts for the injection of material into the network upon fusion through the term $+k_{u}(h_{0}\ell_{0}-h_{c})$, where $k_{u}$ is the rate for each cluster to fuse with the source. This term arises from the fact that whenever a cluster fuses with the source cluster, both clusters leave the interaction with concentration $h_{0}=1/\ell_{0}$, meaning that the material per unit in the non-source cluster increases by ($h_{0}\ell_{0}-h_{c}$). With these modifications, we arrive at the dynamic equation for per-unit concentration within the contact zone:
@@
$\begin{gathered}h(b)=\frac{A}{b}\left(e^{b/\lambda}+\frac{B}{A}e^{-b/\lambda}\right)=h_{c}\\ \rightarrow\quad A=\frac{(R+\lambda)e^{-b/\lambda}bh_{c}}{(R+\lambda)+(R-\lambda)e^{2(R-b)}}\end{gathered}$ (18)

### v2

@@ Results
$h_{c}$ $=\frac{\ell_{0}h_{0}^{(d)}}{1+k_{d}/(fk_{u})+k_{d}z/(fk_{u}v_{c})},$ (4c)
where the first term in the updated $S$ accounts for material in the source cluster and the second term represents material in the rest of the network. When clusters are large and interactions are infrequent, the above solution reduces to the static network limit of Eq. 1. When the clusters are small, the source cluster is fully filled ($h_{0}^{(d)}\rightarrow h_{0}$) and so is each cluster that interacts with it ($f\rightarrow 1$). The general solution then approaches the fragmented social network limit of Eq. 3.
In Fig. 4, we show the spreading of material through simulated networks with increasing connectivity, spanning across the different regimes. In the limit of rapid decay (small $\tau_{\textrm{d}}$), material spreads within a single cluster of dimension $d$, with the amount delivered scaling as $S\sim\tau_{d}^{d/2}$ When particle diffusivity is fast and clusters are small, the source cluster fills faster than interactions can occur and the amount of material plateaus at the size of a single cluster, in the regime of $\tau_{c}<\tau_{d}<\tau_{\text{int}}$. For longer decay times, interaction between distinct clusters dominates the spread, and the total material scales linearly as $S\sim\tau_{d}$. When particle diffusion is slow (Fig. 4b), the cluster containing the source unit is able to offload material through interactions even before it is fully filled ($\tau_{\text{int}}<\tau_{c}$), and the plateau region narrows or disappears.
The simplified analytic model (Eq. 4) approximately matches simulations over a broad range of network connectivities and timescales (Fig. 4). However, some discrepancy occurs for networks that are close to the percolation transition (light green curve), particularly when the particle diffusivity is low. In this case, the approximation systematically overestimates material spreading through the network. This may be the result of transient fissions within a single cluster (which could temporarily hinder material spreading [21]), or steric inaccessibility of nodes buried within a cluster, neither of which are accounted for in our analytic models. Additionally, the percolation transition corresponds to a broad variability in cluster sizes [13], resulting in the source unit occasionally being trapped in very small clusters that limit material delivery.
@@
Spreading rates on mammalian mitochondrial networks
Figure 5: Extracted structures of mammalian mitochondrial networks exemplify both the static network and social network regimes. (a) Representative raw, segmented, and skeletonized images (top to bottom) for mitochondria from SH-SY5Y (human neuroblastoma), IMR90 (human fibroblast), and U2OS (human osteosarcoma) cells. Segmented images show each connected cluster with a different color. Scale bar: $10\mu\text{m}$. (b) Time to fill half of the mitochondrial network with material supplied from a fixed source, as predicted by the analytic model, is plotted as a function of the material diffusivity for each cell type. Solid curves correspond to a ‘prototypical cell’, where parameters are averaged within each cell type before plugging into the analytic model. Dashed lines indicate results for individual cells. (c) The distribution of cells by mean cluster size fraction (mean cluster length divided by total mitochondrial network length). The dashed vertical line indicates the threshold ($0.29$) to classify networks as disconnected or hyperfused, as determined by k-means clustering with k=2. (d) Network filling half-times as in (b), with cells grouped by the mitochondrial network class (disconnected or hyperfused). Solid curves are obtained by averaging half-times across individual cells within each class with shaded regions indicating the standard error of the mean. Individual cell results are shown as dashed lines.
To illustrate the utility of our model in providing directly testable predictions for mitochondrial mixing, we analyze time-resolved images of fluorescently tagged mitochondria in three human cell lines: SH-SY5Y, IMR90, and U2OS. Specifically, we segment, skeletonize, and track the mitochondria of each cell in our data set (Figure 5a and Supplemental Video 2), extracting structural metrics (degree distribution, cluster size, fractal dimension) and dynamic parameters (fission rate and mitochondrial diffusivity). These values (provided in Supplemental Material) are used to parameterize our analytic model, which provides a quantitative estimate of mitochondrial material spreading. We estimate the time required for material to fill half the network ($t_{1/2}$) as the value of $\tau_{d}$ corresponding to $S$ reaching half its maximal value. This metric provides a prediction for how quickly particles of different types should spread through most of the mitochondrial network.
Figure 5b shows our predictions of spreading time on prototypical ‘average cells’, with individual parameters averaged across all the cells of that type. The spreading time is plotted as a function of the material diffusivity, $D_{p}$, in real units. For typical mitochondrial matrix proteins with $D_{p}=20\mu\text{m}^{2}\text{/s}$ [66], we expect a spreading time of $\sim 4$ min for SH-SY5Y, and $\sim 0.5$ hr for IMR90 and U2OS cells. Notably, the different cell types exhibit distinct behaviors. For IMR90 and U2OS cells, the spreading time asymptotes for particle diffusivities above roughly $1\mu\text{m}^{2}/\text{s}$. By contrast, in the typical SH-SY5Y cell, spreading times continue to decrease with higher material diffusivity. These results arise primarily from differences in the network architecture: SH-SY5Y mitochondria tend to be hyperfused, while IMR90 and U2OS exhibit more disconnected networks, so that dispersion becomes limited by interactions between distinct mitochondrial clusters rather than diffusion within connected mitochondria. Overall, our results indicate that cultured human cell lines can vary substantially in their mitochondrial architecture and mixing dynamics.
It should be noted, however, that there is extensive variation in the mitochondrial structure and associated network parameters between individual cells of the same type (dashed lines in Fig. 5b). To further explore the dependence of spreading rates on network connectivity, we reclassified all cells in our data set as ‘disconnected’ or ‘hyperfused’ based on the mean cluster size as a fraction of the total mitochondrial population. Figure 5c shows the distribution of cells by mean cluster size fraction and the classification threshold obtained via a k-means clustering algorthim. All cells with mean cluster size fraction below the threshold were classified as disconnected, with the remainder considered hyperfused. Figure 5d demonstrates the difference in spreading behavior for these two cell classes, with hyperfused cells exhibiting strong dependence on material diffusivity across the entire relevant range. Spreading on disconnected networks plateaus for particles with diffusivities above $D_{p}\approx 1\mu\text{m}^{2}\text{/s}$. We note that for slowly diffusing material, disconnected networks can actually increase the spreading rate. This is due to the mobility of mitochondrial clusters, which allows each unit to interact with more neighbors than they could contact in the static low-dimensional hyperfused networks ($1\leq d\lesssim 2$).
The mitochondrial networks for different cell types exhibit multiple morphological differences. In particular, the total network size in U2OS cells is substantially larger than in the other two cell types (Fig. 5a; see Supplemental Material). However, because the mean cluster size fraction is similar in both IMR90 and U2OS cells, the predicted filling half-time is similar as well. For systems that fall into the disconnected social network regime, it is the relative size of the individual mitochondrial clusters compared to the full network that determines filling times.
The distinction in predicted spreading rates between different cell types can be attributed to their different propensities for mitochondrial connectivity. Many (though not all) SH-SY5Y cells exhibit hyperfused networks, while IMR90 and U2OS cells fall primarily in the disconnected category. SH-SY5Y cells thus serve as examples of the stationary physical network regime, where spreading is limited by material transport along a fractal-dimensional continuum. The other two cell types have mitochondria that behave akin to a social network, with interactions between clusters limiting the spread of rapidly diffusing particles. Notably, for both IMR90 and U2OS mitochondrial networks, the spreading at high $D_{p}$ is limited by mitochondrial mobility rather than the timescale of local fusion or fission (see Supplemental Appendix). We note that the balance of fusion to fission rates, which sets network connectivity, plays an important role in dictating material spread. However, our analysis indicates that the interaction rate is limited by mitochondria moving through space to encounter new fusion partners rather than by the waiting time for fusion between two nearby mitochondria ($\tau_{\text{enc}}\gg\tau_{u}$). Thus, the model predicts that proportional increases in fusion and fission rates should have little effect on mitochondrial mixing.
@@ Discussion
In this work, we develop a mathematical framework for mixing of diffusive material in dynamic networks and apply it to mitochondrial networks in cultured human cells. We examine two limiting regimes: one of static, well-connected physical networks and one of socially interacting homogeneous clusters. In physical networks, spreading rates increase with material diffusivity and network dimensionality. In three-dimensional social networks, material accumulates linearly in time and is limited either by the fusion rate between nearby clusters or the mobility of those clusters.
@@
The models presented here are based on measurable structural and dynamic features. The static model requires network dimensionality, cluster size, and particle diffusivity, while the social model additionally needs the fusion rate and cluster mobility as inputs. Modern imaging techniques allow quantification of mitochondrial network structure and dynamics in a variety of cellular systems  [74, 28, 75, 13, 32]. Our results connect these morphological measurements to the dynamics of material spreading through the mitochondrial population.
We show example calculations for three human cell types, demonstrating the distinct parameter regimes accessed by their mitochondrial networks. For the set of cells considered here, many SH-SY5Y cells fall in the well-connected regime, while the majority of IMR90 and U2OS cells exhibit disconnected social mitochondrial networks, whose mixing is limited by mitochondrial mobility and encounter. It should be noted, however, that the mitochondrial structure varies substantially among individual cells of a given type, and is also expected to change as a function of metabolic conditions [38] and genetic perturbations [75, 37]. The predictions made by this model could be tested in future work by quantifying the dispersion of locally photoconverted proteins through mitochondrial networks of different architectures [38].
@@
Our quantitative calculations focus on a simplified dynamic system, with a single source of material held at fixed concentration. Realistic biological scenarios are of course likely to involve additional complications, including multiple sources (such as the mitochondrial nucleoids), non-trivial regulation at the source, or diffusive barriers within the mitochondria [63]. We also assumed that individual mitochondrial units are all identical in their fusion and fission behavior, whereas it is possible that some units remain isolated from the rest of the network, while others are more likely to engage in interactions. Furthermore, dynamic rearrangements beyond fusion and fission (such as tubule extension and branch sliding) may contribute to mixing within individual mitochondrial network components [76, 77, 78]. However, the minimal model presented here provides a basic building block from which more complicated scenarios can be constructed and which can serve as a null hypothesis for analyzing experimental observations.
This work brings us closer to understanding the functional implications of variable network connectivity for the spread of ions, lipids, proteins, and genetic information within mitochondria. For example, diffusive spreading dictates the heterogeneity of mitochondrial contents arising from mRNA and protein import site localization on the mitochondrial surface [79, 58]. Similarly, diffusion of mitochondrial transcription factors and mRNAs may dictate the spheres of influence of individual nucleoids within the network [80].
Mitochondrial networks are known to alter their architecture in response to changing environmental conditions, with excess glucose causing network fragmentation while starvation results in hyperfused networks [38]. A quantitative framework for material spreading in different structures provides insight on the implications of these transitions for many biologically critical processes, such as the distribution of harmful reactive oxygen species [47], the ability to isolate or complement deleterious DNA mutations [81, 80], and the tunneling of calcium ions, proton gradients, and ATP throughout the cell [82, 48, 49]
@@
While we have focused on mitochondria, analogous problems of diffusive spreading arise in many systems, such as porous media in the geosciences [23], trophallactic networks in honeybees [4], and Alzheimers progression across connected regions in the brain [83]. By bringing together models of static physical networks and transiently interacting social networks, our results enable a quantitatively predictive link from network architecture to the rate of diffusive dispersion.
@@ Acknowledgements.
EFK acknowledges funding from NSF grant PHYS-2310229, the Chan Zuckerberg Initiative, and a grant from the UCSD Academic Senate. SL acknowledges funding by CZI Metabolism Across Scales grant #RR-8817, and by NSF CAREER award #2339182. We are grateful to Austin EYT Lefebvre, Gabriel Sturm, Nigel Goldenfeld, and Suliana Manley and to the members of the CZI Theory Institute Without Walls for helpful discussions.
@@
$\begin{split}S=\frac{\lambda_{p}}{\ell_{0}}\left[\tanh\left(\frac{R_{n}}{2\lambda_{p}}\right)+\tanh\left(\frac{3R_{n}}{2\lambda_{p}}\right)\right]\end{split}$ (7)
@@
An additional adjustment accounts for the injection of material into the network upon fusion through the term $k_{u}(h_{0}\ell_{0}-h_{c})$, where $k_{u}$ is the rate for each cluster to fuse with the source. This term arises from the fact that whenever a cluster fuses with the source cluster, both clusters leave the interaction with concentration $h_{0}=1/\ell_{0}$, meaning that the material per unit in the non-source cluster increases by ($h_{0}\ell_{0}-h_{c}$). With these modifications, we arrive at the dynamic equation for per-unit concentration within the contact zone:
@@
$\begin{gathered}h(b)=\frac{A}{b}\left(e^{b/\lambda}+\frac{B}{A}e^{-b/\lambda}\right)=h_{c}\\ \rightarrow\quad A=\frac{(R+\lambda)e^{-b/\lambda}bh_{c}}{(R+\lambda)+(R-\lambda)e^{2(R-b)/\lambda}}\end{gathered}$ (18)
@@
.6 Experimental Imaging and Data Processing
.6.1 Mammalian cell growth, transfection, and vital dyes
Human IMR90 cells (American Type Culture Collection #CCL-186) and U2OS-Cas9 cells (HD Cas9-012) were grown in high-glucose Dulbecco’s modified Eagle’s medium (DMEM) supplemented with 10% heat-inactivated fetal bovine serum, 1% L-Glutamine (100X), and 1% penicillin/streptomycin at 37C in a humidified 5% CO2 chamber. Human SH-SY5Y cells (American Type Culture Collection #CRL-2266) were grown in high glucose Dulbecco’s modified Eagle’s medium (DMEM) supplemented with 10% heat-inactivated fetal bovine serum, 1% non essential amino acids (NEAA) and 1% Sodium Pyruvate (NaPyr) at 37C in a humidified 5% CO2 chamber. Before imaging, cells were seeded onto glass-bottom 35-mm dishes (Mattek) and cultured for 24 to 48 hours. Transient transfection was performed using Lipofectamine 2000 according to the manufacturer’s protocol (Thermo Fisher Scientific) and cells were imaged 24 hours later, unless otherwise noted. U2OS-Cas9 cells were transfected with non-targeting Scrambled sgRNA#1 (EditCo) using RNAiMAX Transfection Reagent according to the manufacturer’s protocol (Thermo Fisher Scientific), media was replaced 24 hours later, and cells imaged 5 days later. For live cell imaging, 1 ml of conditioned cell media was removed and reserved prior to staining with 50 nM MitoTracker Deep Red (Thermo Fisher Scientific) directly in the imaging dish for 30 min. Medium was then replaced with the reserved conditioned media just before imaging. Plasmids and resource identifiers are reported in Table 1.
Plasmid Source and Reference
Mito-BFP mito-BFP was a gift from Gia Voeltz (Addgene plasmid # 49151; http://n2t.net/addgene:49151; RRID:Addgene_49151). https://pubmed.ncbi.nlm.nih.gov/21885730/
Su9-EGFP Su9-EGFP was a gift from David Chan (Addgene plasmid # 23214; http://n2t.net/addgene:23214; RRID:Addgene_23214). https://pubmed.ncbi.nlm.nih.gov/12527753/
Table 1: Plasmids
.6.2 Microscopy and image acquisition
All images were acquired using a Zeiss LSM 980 with Airyscan 2 laser scanning confocal microscope, equipped with 405-nm, 488-nm, 561-nm, and 639-nm laser lines and Fast Airyscan detector array. Images were acquired using an inverted 63x/1.4 numerical aperture oil objective. All live imaging was done in a humidified chamber at 37°C and in the presence of 5% CO2. Airyscan processing was performed using Zeiss ZEN Blue software version 3.7 (Carl Zeiss). Image brightness and/or contrast were linearly adjusted in ZEN Blue.
.6.3 Network Extraction
Individual cells are separated and manually cropped in ImageJ using a max projection in Z and time. Areas and volumes are estimated by measuring the cell area and subtracting the nuclear area, with height assumed uniform throughout. The mitochondria for each cell are segmented and tracked using the software package Nellie  [87]. Skeletons and corresponding network structures are extracted from the Nellie segmentation at each time point using the skimage [88] and skan [89] python packages. Degree distributions for the network are computed as in [57] to provide for an assumed underlying mitochondrial unit length of $0.5\mu\text{m}$, matching simulations. We then extract all structural parameters (cluster sizes, radius of gyration, fractal dimension, mean graph distance, etc.) for each imaged network in the same way as for simulated networks, using a custom network analysis package in Matlab [90].
.6.4 Dynamic measurements
There are two dynamic parameters which must be extracted from experimental data in order to apply the analytic model. From Nellie outputs of segmented and time-linked objects, we extract the trajectories of individual labeled mitochondria, as marked by their center-of-mass positions in consecutive frames. Trajectories containing fewer than three time points or very fast motion of greater than $0.64\mu\text{m/s}$ are discarded. We restrict our analysis to mitochondria of length $2\mu\text{m}$ or less, as large clusters are more prone to tracking artifacts. Positions are projected into the xy-plane and mean squared displacements (MSD) are calculated between consecutive frames on each trajectory. Finally, the effective mitochondrial diffusivity for a given cell is estimated as: $D_{\text{mito}}=\langle\Delta r^{2}\rangle/(4\Delta t)$, where $\langle\Delta r^{2}\rangle$ is the MSD between consecutive frames and $\Delta t$ is the time between frames.
Second, fission events are manually identified in ImageJ using the max projection in Z for each cell. To calculate the effective per-node fission rate constant, $k_{f}$, we then divide these fission counts by the elapsed time for the movie and a weighted sum of junction nodes ($x_{2}+\frac{3}{2}x_{3}$), with $x_{i}$ the average number of degree-$i$ nodes in the network [13].
It is important to make a distinction between microscopic fusion and fission events, which occur with local rate constants $k_{f},k_{u}$, and large-scale events that encompass network rearrangement. The latter include ‘novel fusions’ (fusion with a new network unit) and ‘complete fission’ events that proceed them. As found in [57], for simulated mitochondrial networks, the complete fission rate constant is roughly ten-fold lower than the local fission rate constant $k_{f}$. Because microscopic fissions are quickly followed by re-fusion and are unlikely to be observed in our time-lapse images, we assume these counts correspond to complete fission events only, and set $k_{f}$ to be ten-fold higher than the observed rate. Finally, the cluster-cluster fusion rate $k_{u}$ is estimated as $k_{u}\approx\langle k_{f}\rangle(C_{1}/\delta v_{1}+C_{2}/\delta v_{2})$ where $C_{1}=V_{\text{cell}}x_{2}/(\frac{1}{2}x_{1}^{2})$, $C_{2}=\frac{3}{2}V_{\text{cell}}x_{3}/(x_{1}x_{2})$, and $\delta v_{1},\delta v_{2}$ are the contact volumes per node for tip-tip and tip-side fusion [57, 13].
We note that our estimates for the local fission rate are likely an underestimate, as we are limited by framerate and the ability to manually identify all fission events within a given movie. Estimates for the mitochondrial diffusivity, on the other hand, are likely a slight overestimate, as tracking artifacts may cause artificial jumps in trajectories which cannot be completely filtered out. However, these limitations do not alter the results plotted in Fig. 5b,d because the mitochondria in these cells are already in a mobility-limited regime. This means that when spreading occurs via cluster-cluster interactions as described by the social network model, the interaction time is dominated by the timescale to encounter new clusters, $\tau_{\text{enc}}$.
We can see quantitatively that the mitochondria in our data set fall into this regime by analyzing the ratio: $\tau_{\text{enc}}/\tau_{u}=\rho v_{c}k_{u}/(\rho 4\pi Db)$, which compares the total fusion rate in the contact zone to the rate of mitochondria arriving at the contact zone. For the cells in our data set, this ratio ranges from $10^{1}-10^{4}$, with the majority of cells centered around $10^{2}$. Because our estimate for $k_{u}$ is a lower bound and our estimate for $D$ is an upper bound, the predicted ratio of $\tau_{\text{enc}}/\tau_{u}$ must also be a lower bound. Therefore we can safely say that the results for spreading in the mitochondrial networks used for this analysis are insensitive to potential inaccuracy in the measured fission rate.
.6.5 Solving the social network model in a slab-like geometry
As the cells we are analyzing are nearly flat (area $\sim 1000\mu\text{m}^{2}$, height $\sim 2\mu\text{m}$), we solved the social model (Eq. 2) in a quasi-2D slablike geometry to make the predictions shown in Fig. 5. This amounts to solving the diffusion equation in cylindrical coordinates, with a result that looks exactly the same as Eq. 3 up to a redefinition of the spatial spreading term, contact volume, and mitochondrial unit density:
$\frac{z}{2\pi bH\lambda}=$ $\frac{-I_{1}(b/\lambda)K_{1}(R/\lambda)+K_{1}(b/\lambda)I_{1}(R/\lambda)}{I_{0}(b/\lambda)K_{1}(R/\lambda)+K_{0}(b/\lambda)I_{1}(R/\lambda)},$ (24a)
$v_{c}=$ $\pi(b^{2}-a^{2})H,$ (24b)
$\rho=$ $(N_{0}-\langle n\rangle)/[\pi(R^{2}-a^{2})H],$ (24c)
where $I_{j},K_{j}$ are modified Bessel functions, $H$ is the height of the cell and we calculate $R=\sqrt{A_{\text{cell}}/\pi}$, the effective planar radius of the cell based on the measured cell area. The steric radius $a$ is set based on the radius of gyration averaged over all clusters and $b$ is calculated from Eq. 24b, analogously to the 3D simulations.
In 4 of the analyzed SH-SY5Y cells, the average mitochondrial cluster size was greater than half of the total mitochondrial content in the cell. As the social network model assumes multiple interacting clusters of similar size, we used only the stationary network approximation to estimate the spreading times for these networks, with the source assumed to be located on a cluster of the typical largest size (for these networks, $\sim 0.75$ of the total mitochondrial content).
Supplemental Video Caption
Supplemental Video 1: Steady-state dynamics of the mitochondrial network model. Material concentration is indicated on a white-to-red gradient. A single fixed-concentration unit serves as the source. Top: disconnected network, Middle: network near percolation transition, Bottom: hyperfused network. Simulation parameters correspond to those listed in Fig. 1b. Each movie runs over a simulated time of $60$ dimensionless units (corresponding to $\approx 120$min).
Supplemental Video 2: Example live-cell imaging data of mitochondrial networks in SH-SY5Y (top), IMR90 (middle), and U2OS (bottom) cells. Raw (white) and segmented (multicolored) mitochondria are shown, with all segmentations obtained using Nellie [87]. Connected clusters are indicated by different colors in the initial frame, with subsequent voxel colors determined via Nellie’s flow mapping algorithm. All movies run backward then forward over a span of 116 seconds.